# Control the qubit-qubit coupling with double superconducting resonators

### v1

@@ Abstract
We experimentally studied the switching off processes in the double-resonator coupler superconducting quantum circuit. In both frequency and time-domain, we observed the variation of qubit-qubit effective coupling by tuning qubits’ frequencies. According to the measurement results, by just shifting qubits’ frequencies smaller than 50 MHz, the effective qubit-qubit coupling strength can be tuned from switching off point to two qubit gate point (effective coupling larger than 5 MHz) in double-resonator superconducting quantum circuit. The double-resonator coupler superconducting quantum circuit has the advantage of simple fabrications, introducing less flux noises, reducing occupancy of dilution refrigerator cables, which might supply a promising platform for future large-scale superconducting quantum processors.
@@ I Introduction
In this article, we experimentally studied the switching off processes in the double-resonator coupler superconducting quantum circuit. In both frequency and time-domain, we measured the variation of qubit-qubit coupling strength by tuning frequency differences of qubits relative to the resonators coupler through the Z-control signals. By changing the energy level crossing points relative to the fixe frequency resonator with the Z-control signals (DC-biased current), we observed the variation of qubit-qubit anti-crossing gap with two-tone spectroscopy on superconducting quantum circuits. Even with the relative noise measurement data in time-domain measurement(without Josephson parameter amplifier and high base-temperature of dilution refrigerator above 25mK), we still can see the variation of envelope of vacuum Rabi oscillation which reflects the change of effective qubit-qubit coupling. According to the measurement results, the effective qubit-qubit coupling can be tuned from switching off point to two qubit gate point (effective coupling larger than 5 MHz) by just shifting about 50 MHz of qubits’ frequencies in double-resonator superconducting quantum circuit.
@@ II Theoretical model
If the transition frequencies of two qubits are tuned between that of two resonator coupler ($\omega_{b}>\omega_{1,2}>\omega_{a}$), the high and low frequency resonators will respectively make a negative and positive contributions to the qubit-qubit interaction. When the net contribution of two resonators is same value but opposite sign with the direct qubit-qubit coupling $g_{12}$, the qubit-qubit coupling can be switched off. Even for extreme small (or even zero) value of $g_{12}$, the switching off ($g^{(eff)}_{12}=0$) can be realized without requiring large-scale of frequency shift in the double-resonator coupler superconducting quantum circuit.
@@
The transition frequency of qubit-2 is fixed at the certain value by DC-bias current, then the effective qubit-qubit coupling depends on the frequency of qubit-1 (or $\Delta_{\lambda\beta}$) according to Eq.(2). If we apply an independent bias current to sweep the frequency of qubit-1 with DC-biased frequency around transition frequency of qubit-2, then the variation of qubit-qubit anti-crossing gap are shown in Fig.3. By fixing the frequency of qubit-2 at different frequency with DC bias current, and the frequency of qubit-1 sweep around the frequency of qubit-2. The anti-gap reduce from about 10 MHz (4.58 GHz) in Fig.3(a) and reduce to below 2 MHz in Figs.3(d) and 3(e) (close to 4.62 GHz). If qubit-2 is tuned to about 4.37 GHz, the anti-crossing gap is almost invisible in Fig.3(e) where the qubit-qubit coupling is turned off. By choosing direct qubit-qubit coupling as 0.88 MHz, with the qubit-resonator coupling strength obtained from the anti-crossing gap, the calculated switching off point coincide well the measurement results in Fig.(3).
@@
By only shifting 50 MHz from the switching off points with flux pulse, the anti-crossing gap effective qubit-qubit coupling can be tuned from 0 Hz to about 5 MHz as shown in Fig.3(a). Thus the switching off positions can be very close to two-qubit gate regimes (about 5 MHz effective qubit-qubit coupling), thus the flux noises can be greatly suppressed during the quantum operations for the double-resonator couplers superconducting quantum chip. The double-resonators coupler can turn off the qubit-qbuit coupling without requiring additional flux lines, this can reduce flux noises and the occupation number for high-frequency cables of dilution refrigerator.
@@ IV Time-domain measurement
In the absence of Josephson parameter amplifier and the relative high base temperature (above 25mK), the Signal-to-Noise Ratio is not high in this measurement. If we use the normalization method to calculate excited state occupation probabilities, more errors will be induced by the noise data. As shown in Figs.6(b)-6(h), we use $\Delta|IQ|(=|IQ|-baseline)$ to describe the variation of the readout signal which can reflect envelope variations of vacuum Rabi oscillation. Even with low SNR, we still can see some feature of envelope variation of $\Delta|IQ|$ from the noisy data (see Figs.6(b)-6(h)). When the flux pulse amplitude is close to -6.0 mV, the envelope of $\Delta|IQ|$ is not clear in Fig.6(e), which should be close to switching off point(see fig.3). As the change of the flux pulse amplitude of qubit-1, the envelope for the curve of $\Delta|IQ|$ become clearer at larger frequency detuning ($|\Delta_{12}|$) because of nonzero energy exchange between two qubits, the oscillation periods of $\Delta|IQ|$ reflect variation of qubit-qubit coupling strength. The switching off point measured by the Vacuum Rabi oscillation roughly coincide with the anti-crossing gap in the frequency domain. The measurement result should be enhanced in future experiment if we add the Josephosn parameter amplifier, Purcell filter, and enhance the effective qubit-qubit coupling.
@@ V Conclusions
In conclusion, we experimentally studied tunable qubit-qubit coupling in the double-resonator coupler type superconducting quantum circuit. The induced indirect qubit-qubit coupling by two resonators can be cancelled, so the switching off can be realized without the direct qubit-qubit coupling. In the frequency domain and time-domain, we observed the variation of the effective qubit-qubit induced by the two-fixed frequency resonators. In future experiment, we will add the Josephson parameter amplifier, Purcell filter, and also conduct the calibration of Z-pulse distortion to improve the measurement result.
@@ VI ACKNOWLEDGMENTS
Hui Wang thank Zhiguang Yan for valuable suggestions on the hardware setup, measurement codes writing, and measurement result analysis. We also thank Shiyu Wang and Russell S. Deacon for friendly help during the measurement and fabrication of the samples, and also Chih-Yao Shih, Ching-Yeh Chen, Hiroto Mukai, and Hang Xue for maintaining the running of dilution refrigerator. J.S. Tsai is supported by the Japan Science and Technology Agency (Moonshot R D, JPMJMS2067; CREST, JPMJCR1676) and the New Energy and Industrial Technology Development Organization (NEDO, JPNP16007).
@@ VI.1 Experiment setup
The measurement is use the keysight M3202A and digitizer M3102A, the microwave signals are created by the RS sgs100a. The attenuation in the read-in and XY control line are about 60db, and about 44db for the Z-control line The DC-current and the flux pulse is combined by the bias-Tee in room temperature. The infrared filter are installed for the driving and readout cable. The DC current will bias the qubits at suitable frequency points for the Rabi or the vacuum Rabi measurement, while the flux pulse is for the fast sweeping of qubits’ frequency, and they combined together with each other in room temperature with a bias-Tee.

### v2

@@ Abstract
We experimentally studied the switching off processes in the double-resonator coupler superconducting quantum circuit. In both frequency and time-domain, we observed the variation of qubit-qubit effective coupling by tuning the frequency differences between qubits and the double-resonator coupler. According to the measurement results, by just shifting about 50 MHz of qubits’ frequencies, we can tune the effective qubit-qubit coupling strength from switching off point to two qubit gate point (effective coupling larger than 5 MHz) in double-resonator superconducting quantum circuit. The double-resonator (coupler) superconducting quantum circuit has the advantage of simple fabrications, introducing less flux noises, reducing occupancy of dilution refrigerator cables, which might supply a promising platform for future large-scale superconducting quantum processors.
@@ I Introduction
In this article, we experimentally studied the switching off processes in the double-resonator coupler superconducting quantum circuit. In both frequency and time-domain, we measured the variation of qubit-qubit coupling strength by tuning frequency differences of qubits relative to the resonators coupler through the Z-control signals. By changing the energy level crossing points relative to the fixe frequency resonator with the Z-control signals (DC-bias current), we observed the variation of qubit-qubit anti-crossing gap with two-tone spectroscopy on superconducting quantum circuits. Even with the relative noise measurement data in time-domain measurement(without Josephson parameter amplifier and high base-temperature of dilution refrigerator above 25mK), we still can see the variation of envelope of vacuum Rabi oscillation which reflects the change of effective qubit-qubit coupling. According to the measurement results, the effective qubit-qubit coupling can be tuned from switching off point to two qubit gate point (effective coupling larger than 5 MHz) by just shifting about 50 MHz of qubits’ frequencies in double-resonator superconducting quantum circuit.
@@ II Theoretical model
If the transition frequencies of two qubits are tuned between that of two resonator coupler ($\omega_{b}>\omega_{1,2}>\omega_{a}$), the high and low frequency resonators will respectively make a negative and positive contributions to the qubit-qubit interaction. When the net contribution of two resonators is same value but opposite sign with the direct qubit-qubit coupling $g_{12}$, the qubit-qubit coupling can be switched off. Even for extreme small (or even zero) value of $g_{12}$, the switching off ($g_{eff}=0$) can be realized without requiring large-scale of frequency shift in the double-resonator coupler superconducting quantum circuit.
@@
The transition frequency of qubit–2 is fixed at the certain value by DC-bias current, then the effective qubit-qubit coupling depends on the frequency of qubit-1 (or $\Delta_{\lambda\beta}$) according to Eq.(2). If we apply an independent bias current to sweep the frequency of qubit-1 with DC-biased frequency around transition frequency of qubit-2, then the variation of qubit-qubit anti-crossing gap are shown in Fig.3. By fixing the frequency of qubit-2 at different frequency with DC bias current, and the frequency of qubit-1 sweep around the frequency of qubit-2. The anti-gap reduce from about 10 MHz (4.58 GHz) in Fig.3(a) and reduce to below 2 MHz in Figs.3(d) and 3(e) (close to 4.62 GHz). If qubit-2 is tuned to about 4.37 GHz, the anti-crossing gap is almost invisible in Fig.3(e) where the qubit-qubit coupling is turned off. By choosing direct qubit-qubit coupling as 0.88 MHz, with the qubit-resonator coupling strength obtained from the anti-crossing gap, the calculated switching off point coincide well the measurement results in Fig.(3).
@@
By only shifting 50 MHz from the switching off points with flux pulse, the anti-crossing gap effective qubit-qubit coupling can be tuned from 0 Hz to about 5 MHz as shown in Fig.3(a). Thus the switching off positions can be very close to two-qubit gate regimes (about 5 MHz effective qubit-qubit coupling), thus the flux noises can be greatly suppressed during the quantum operations for the double-resonator couplers superconducting quantum chip. The double-resonators coupler can turn off the qubit-qbuit coupling without requiring additional flux lines, this can reduce flux noises and the occupation number for high-frequency cables of dilution refrigerator. The width of coupler resonator can be narrower in future design, than the size of resonator couplers can be much smaller space of the chip. To show the value of qubit-resonator strength, the maximal frequency of one qubit is larger than the high frequency resonator-coupler which is not necessary in the real superconducting quantum processor.
@@ IV Time-domain measurement
In the absence of Josephson parameter amplifier and the relative high base temperature (above 25mK), the Signal-to-Noise Ratio is not high in this measurement. If we use the normalization method to calculate excited state occupation probabilities, more errors will be induced by the noise data. As shown in Figs.6(b)-6(h), we use $\Delta|IQ|(=|IQ|-baseline)$ to describe the variation of the readout signal which can reflect envelope variations of vacuum Rabi oscillation. Even with low SNR, we still can see some feature of envelope variation of $\Delta|IQ|$ from the noisy data (see Figs.6(b)-6(h)). When the flux pulse amplitude is close to -6.0 mV, the envelope of $\Delta|IQ|$ is not clear in Fig.6(e), which should be close to switching off point(see fig.3). As the change of the flux pulse amplitude of qubit-1, the envelope for the curve of $\Delta|IQ|$ become clearer at larger frequency detuning ($|\Delta_{12}|$) because of nonzero energy exchange between two qubits, the oscillation periods of $\Delta|IQ|$ reflect variation of qubit-qubit coupling strength. The switching off point measured by the Vacuum Rabi oscillation roughly coincide with the anti-crossing gap in the frequency domain. The measurement result should be enhanced if the Josephosn parameter amplifier, Purcell filter, Z-pulse distortions are applied in future experiment.
@@ V Conclusions
In conclusion, we experimentally studied tunable qubit-qubit coupling in the double-resonator coupler type superconducting quantum circuit. The induced indirect qubit-qubit coupling by two resonators can be cancelled, so the switching off can be realized without the direct qubit-qubit coupling. In the frequency domain and time-domain, we observed the variation of the effective qubit-qubit coupling tuning by frequency detuning between the qubits and resonator couplers. By reducing the designed line widths of resonator-coupler, than occupation space of resonator couplers can be much smaller in the multi-qubit superconducting quantum chips.
@@ VI ACKNOWLEDGMENTS
Hui Wang thank Zhiguang Yan for valuable suggestions on the hardware setup, codes writing, and result analysis of the low-temperature measurments. We also thank Shiyu Wang and Russell S. Deacon for friendly help during the measurements and fabrications of the superconducting samples. J.S. Tsai is supported by the Japan Science and Technology Agency (Moonshot R D, JPMJMS2067; CREST, JPMJCR1676) and the New Energy and Industrial Technology Development Organization (NEDO, JPNP16007).
@@ VI.1 Experiment setup
The measurement is use the keysight M3202A and digitizer M3102A, the microwave signals are created by the RS sgs100a. The attenuation in the read-in and XY control line are about 60db, and about 44db for the Z-control line The DC-current and the flux pulse is combined by the bias-Tee in room temperature. The infrared filters are installed for the driving and readout cables. The DC current will bias the qubits at suitable frequency points for the Rabi or the vacuum Rabi measurement, while the flux pulse is for the fast sweeping of qubits’ frequency, and they combined together with each other in room temperature with a bias-Tee.